OpenStudy (anonymous):

Indicate the equation of the line that is the perpendicular bisector of the segment with endpoints (4, 1) and (2, -5).

OpenStudy (cwrw238):

m = slope of line = -5-1 / 2-4 = 3 usine y-y1 = m(x-x1) where x1, y1 = 4,1 y - 1 = 3(x - 4) y - 1 = 3x - 12 y - 3x = -11